OpenStudy (anonymous):

Chapter 5:Indices and Logarithms Double checking

OpenStudy (anonymous):

simplify the following

OpenStudy (anonymous):

\[=2^n-2^{n+1}\]

OpenStudy (anonymous):

\[=(2^n-2^n~\times2^{-1})\]

OpenStudy (anonymous):

\[=2^n(1-1-1)\]

OpenStudy (anonymous):

\[=-2^n\]

OpenStudy (butterflydreamer):

correct :)!

OpenStudy (anonymous):

Thnx @butterflydreamer :)

OpenStudy (butterflydreamer):

well your answer is correct but i think .. the working out had an error? :O Idk i solved it differently

OpenStudy (anonymous):

can u show ur working

OpenStudy (butterflydreamer):